OpenStudy (anonymous):

how do you write x^2+y^2=5 in polar coordinates?

OpenStudy (shamil98):

\[\huge x = r \cos \theta , y = r \sin \theta\]

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

how do i get it to be r^2=something?

OpenStudy (shamil98):

\[\huge r = \sqrt{x^2 + y^2}\]

OpenStudy (anonymous):

When converting to polar, x = r cos theta y = r sin theta Plug in x and y (r cos theta) ^ 2 + (r sin theta) ^2 = 5 r^2 * (cos theta)^2 + r^2 * (sin theta) ^2 = 5 factor out r^2 r^2 (cos theta ^2 + sin theta ^ 2) = 5 We see this trig identity cos theta ^2 + sin theta ^2 = 1 so r^2 (1) = 5 thus, r = sqrt(5) Hope this helped.

OpenStudy (anonymous):

It did :) thanks!